# PharmacoMatch: Efficient 3D Pharmacophore Screening through Neural Subgraph Matching

### v1

@@ 6 Conclusion
Code & data availability
We will make our code publicly available upon acceptance via https://github.com/molinfo-vienna/PharmacoMatch and share the curated datasets via (https://zenodo.org/).
Acknowledgements
We thank Roxane Jacob, Nils Morten Kriege, and Christian Permann from the University of Vienna, Klaus-Jürgen Schleifer from BASF SE, and Andreas Bergner from Boehringer-Ingelheim RCV GmbH & Co KG for fruitful discussions and proof-reading of the manuscript. Financial support received for the Christian Doppler Laboratory for Molecular Informatics in the Biosciences by the Austrian Federal Ministry of Labour and Economy, the National Foundation for Research, Technology and Development, the Christian Doppler Research Association, Boehringer-Ingelheim RCV GmbH & Co KG and BASF SE is gratefully acknowledged.
@@ Acknowledgements
Conceptualization: DR, OW. Methodology: DR, OW. Data Curation & Analysis: DR. Code Implementation & Model Training: DR. CDPKit Software & Support: TS. Investigation: DR. Writing (Original Draft): DR. Writing (Review and Editing): DR, OW, TS, TL. Funding Acquisition: TL. Resources: TL. Supervision: OW, TL. All authors have given approval to the final version of the manuscript.
Competing interests
The authors declare no competing interests.
A Appendix
@@ A.1 Dataset curation & statistics
Unlabeled training data was downloaded from the ChEMBL database to represent small molecules with drug-like properties. At the time of data download, the ChEMBL database contained 2,399,743 unique compounds. We constrained the compound category to "small molecules" and enforced adherence to the Lipinsky rule of five (Lipinski et al. 1997), specifically setting violations to "0," resulting in a refined set of 1,348,115 compounds available for download. The molecules were acquired in the form of Simplified Molecular Input Line Entry System (SMILES) (Weininger 1988) strings. Subsequent to data retrieval, we conducted preprocessing using the database cleaning functionalities of the Chemical Data Processing Toolkit (CDPKit) (Seidel 2024). This process involved the removal of solvents and counter ions, adjustment of protonation states to a physiological pH value, and elimination of duplicate structures, where compounds differing only in their stereo configuration were regarded as duplicates. To prevent data leakage, we carefully removed all structures from the training data that would occur in one of the test sets we used for our benchmark experiments. The final set was comprised of 1,221,098 compounds. For each compound within the dataset, a 3D conformation was generated using the CONFORGE (Seidel et al. 2023) conformer generator from the CDPKit, which was successful for 1,220,104 compounds. To enhance batch diversity, we generated only one conformation per compound for contrastive training. Subsequently, 3D pharmacophores were computed for each conformation, with removal of pharmacophores containing less than four pharmacophoric points. The ultimate dataset comprised 1,217,361 distinct pharmacophores.
Figure A1: Pharmacophoric point statistics of the training data. The respective histograms display the total number of pharmacophoric points and the number of points of specific types per pharmacophore in the training data. The complete training dataset contains 1,217,361 distinct pharmacophores.
Figure A1 shows the frequency of pharmacophores with a specific pharmacophoric point count in the training data. On average, a pharmacophore consists of 13 pharmacophoric points, with the largest pharmacophore in the dataset containing 32 points. Pharmacophores with fewer than four points were omitted during data clean-up. Hydrophobic pharmacophoric points and hydrogen bond acceptors are the most prominent, while hydrogen bond donors and aromatics occur less frequently. Ionizable pharmacophoric points and halogen bond donors are comparatively rare.
@@ A.2 Augmentation module
There are two modes for the displacement of pharmacophoric points. Positive pairs were constructed by displacing the pharmacophoric points within the tolerance sphere of the initial pharmacophore. For simplicity, we assumed the same tolerance sphere radius $r_{T}$ across different pharmacophoric types. The coordinate displacement $(\Delta x,\Delta y,\Delta z)$ was created from spherical coordinates $\phi\sim\mathcal{U}(0,2\pi)$ and $\cos{\theta}\sim\mathcal{U}(-1,1)$, which were drawn at random from a uniform distribution. The coordinate displacement was calculated as
@@
where $\Delta r=r_{T}\sqrt[3]{u}$ and $u\sim\mathcal{U}(0,1)$. Negative pairs were created by displacement of the nodes at the border of the tolerance sphere. This was achieved by random sampling from a sphere surface, i.e. with $u=1$.
@@ A.3 Message passing neural network
where $\mathbf{h}_{i}^{(k)}\in\mathbb{R}^{F^{\prime}}$ denotes the node features of node $i$ at layer $k$, $\mathbf{h}_{i}^{(k-1)}\in\mathbb{R}^{F}$ denotes the node features of node $i$ at layer $k-1$, $\mathbf{e}_{ij}\in\mathbb{R}^{D}$ the edge features of the edge from node $i$ to node $j$, $\gamma^{(k)}$ and $\phi^{(k)}$ are parameterized, differentiable functions, and $\bigoplus$ is an aggregation operator like, e. g., the summation operator (Fey and Lenssen 2019). In our encoder architecture, we employed the following edge-conditioned convolution operator, which was proposed both by Gilmer et al. (2017); Simonovsky and Komodakis (2017):
@@
where $\Theta\in\mathbb{R}^{F\times F{{}^{\prime}}}$ denotes learnable weights and $\psi_{\Theta}(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}^{F\times F{{}^{\prime}}}$ denotes a neural network, in our case an MLP with one hidden layer. These transformations map node features $\mathbf{h}$ into a latent representation that combines pharmacophoric types with distance encodings.
@@ A.4 Encoder implementation
where $\mathbf{h}^{l^{\prime}}_{i}$ represents the latent representation after activation. Updating the feature matrix $l$ times yields the final node representations of the pharmacophoric points.
@@
Figure A2: Architecture of the GNN encoder model.
@@ A.5 Model implementation and training
The GNN was implemented in Python 3.10 with PyTorch (v2.0.1) and the PyTorch Geometric library (v2.3.1) (Fey and Lenssen 2019). Both, model and dataset, were implemented within the PyTorch Lightning (Falcon and The PyTorch Lightning team 2019) framework (v2.1.0). Model training was monitored with Tensorboard (v2.13.0). CDPKit (v1.1.1) was employed for chemical data processing. Software was installed and executed on a Rocky Linux (v9.4) system with x86-64 architecture.
@@
Training was performed on a single NVIDIA GeForce 3090 RTX graphics unit with 24 GB GDDR6X. Training runs were performed for a maximum of 500 epochs with a batch size of 256 pharmacophore graphs. Curriculum learning was applied by gradual enrichment of the dataset with increasingly larger pharmacophore graphs. At training start, only pharmacophore graphs with 4 nodes were considered. After 10 subsequent epochs without considerable minimization of the loss function, pharmacophore graphs with one additional node were added to the training data. The loss function was minimized with the Adam optimizer, we further applied gradient clipping. A training run on the full dataset took approximately 48 hours with the above hardware specifications.
Hyperparameter tuning & model selection
Hyperparameters were optimized through random parameter selection, the tested ranges are summarized in Table A2. Unlabeled data was split into training and validation data with a 98:2 ratio. Training runs were compared using the AUROC value on the validation data. This was calculated by treating the positive and negative pairs as binary labels, and the predictions were based on their respective order embedding penalty, which was calculated with Equation (3). Hyperparameter optimization was performed on a reduced dataset with 100,000 graphs, which took approximately 5 hours per run. The best performing models were retrained on the full dataset. The hyperparameters of the final encoder model are summarized in Table A1. After model selection, the final model performance was tested on virtual screening datasets.
Table A1: Hyperparameters of the best performing encoder model
@@
Table A2: Tested hyperparameter ranges for model training.
Hyperparameter
@@ A.6 Virtual screening
General information about the DUD-E targets is summarized in Table A3. For each target we downloaded the receptor structure from the PDB and created the corresponding interaction pharmacophore with the CDPKit. Vector features were converted into undirected pharmacophoric points with LigandScout (Wolber and Langer 2005). The resulting pharmacophore queries (Figure A3) were used in our virtual screening experiments.
Table A3: DUD-E targets that were selected for bechmarking experiments in this study.
@@
Figure A3: Structure-based pharmacophore queries of ten targets of the DUD-E benchmark dataset.
@@
where $S_{MFP}:\mathcal{P}\times\mathcal{P}\rightarrow\mathbb{Z}_{+}$ counts the number of matched feature pairs and $S_{Geom}:\mathcal{P}\times\mathcal{P}\rightarrow[0,1)$ evaluates their geometric fit.
@@
We measured alignment runtimes using the psdscreen tool from the CDPKit with 128 threads on an AMD EPYC 7713 64-Core Processor, while embedding and matching runtimes with PharmacoMatch were recorded using an NVIDIA GeForce RTX 3090 GPU with 24 GB GDDR6X. Runtime per pharmacophore was estimated by dividing the total runtime by the number of pharmacophores in each dataset, with the final estimate taken as the mean of ten runs. The results report the mean and standard deviation of these estimates across all ten datasets.
ROC curves
The performance metrics of our virtual screening experiments are derived from the ROC curves presented in Figures A4 and Figure A5.
Figure A4: Performance comparison of PharmacoMatch and the alignment algorithm. The ROC-curves display the agreement of the hitlist ranking of the two algorithms for ten targets of the DUD-E benchmark dataset.
Figure A5: Absolute screening performance of PharmacoMatch and the alignment algorithm performance for ten targets of the DUD-E benchmark dataset. The pharmacophore queries were generated from the respective PDB ligand-receptor structures.
@@ A.7 Embedding space visualization
Figure A6: UMAP visualization of the vector embeddings of the ACES target.
Figure A7: UMAP visualization of the vector embeddings of the ANDR target.
Figure A8: UMAP visualization of the vector embeddings of the EGFR target.
Figure A9: UMAP visualization of the vector embeddings of the FA10 target.
Figure A10: UMAP visualization of the vector embeddings of the KIT target.
Figure A11: UMAP visualization of the vector embeddings of the PLK1 target.
Figure A12: UMAP visualization of the vector embeddings of the SRC target.
Figure A13: UMAP visualization of the vector embeddings of the THRB target.
Figure A14: UMAP visualization of the vector embeddings of the UROK target.

### v2

@@ 6 Conclusion
Reproducibility statement
The source code of this project can be found under the following link: https://github.com/molinfo-vienna/PharmacoMatch. Please follow the instructions in the repository for reproduction of our results. Training and test data can be downloaded here: https://doi.org/10.6084/m9.figshare.27061081.
Acknowledgement
We thank Roxane Jacob, Nils Morten Kriege, and Christian Permann from the University of Vienna, Klaus-Jürgen Schleifer from BASF SE, and Andreas Bergner from Boehringer-Ingelheim RCV GmbH & Co KG for fruitful discussions and proof-reading of the manuscript. We also thank the anonymous reviewers for their suggestions and comments. Financial support received for the Christian Doppler Laboratory for Molecular Informatics in the Biosciences by the Austrian Federal Ministry of Labour and Economy, the National Foundation for Research, Technology and Development, the Christian Doppler Research Association, Boehringer-Ingelheim RCV GmbH & Co KG and BASF SE is gratefully acknowledged.
@@ Acknowledgement
Conceptualization: DR, OW. Methodology: DR, OW. Data curation & analysis: DR. Code implementation & model training: DR. CDPKit software & support: TS. Investigation: DR. Writing (original draft): DR. Writing (review and editing): DR, OW, TS, TL. Rebuttal: DR. Funding acquisition: TL. Resources: TL. Supervision: OW, TL. All authors have given approval to the final version of the manuscript.
Appendix A Appendix
@@ A.1 Dataset curation & statistics
Unlabeled training data was downloaded from the ChEMBL database to represent small molecules with drug-like properties. At the time of data download, the ChEMBL database contained 2,399,743 unique compounds. We constrained the compound category to ”small molecules” and enforced adherence to the Lipinsky rule of five (Lipinski et al., 1997), specifically setting violations to ”0,” resulting in a refined set of 1,348,115 compounds available for download. The molecules were acquired in the form of Simplified Molecular Input Line Entry System (SMILES) (Weininger, 1988) strings. Subsequent to data retrieval, we conducted preprocessing using the database cleaning functionalities of the Chemical Data Processing Toolkit (CDPKit) (Seidel, 2024). This process involved the removal of solvents and counter ions, adjustment of protonation states to a physiological pH value, and elimination of duplicate structures, where compounds differing only in their stereo configuration were regarded as duplicates. To prevent data leakage, we carefully removed all structures from the training data that would occur in one of the test sets we used for our benchmark experiments. The final set was comprised of 1,221,098 compounds. For each compound within the dataset, a 3D conformation was generated using the CONFORGE (Seidel et al., 2023) conformer generator from the CDPKit, which was successful for 1,220,104 compounds. To enhance batch diversity, we generated only one conformation per compound for contrastive training. Subsequently, 3D pharmacophores were computed for each conformation, with removal of pharmacophores containing less than four pharmacophoric feature points. The ultimate dataset comprised 1,217,361 distinct pharmacophores. Figure 5 shows the frequency of pharmacophores with a specific pharmacophoric feature point count in the training data. On average, a pharmacophore consists of 12.4 pharmacophoric feature points, with the largest pharmacophore in the dataset containing 32 points. Hydrophobic feature points and hydrogen bond acceptors are the most prominent, while hydrogen bond donors and aromatics occur less frequently. Ionizable feature points and halogen bond donors are comparatively rare.
Figure 5: Pharmacophoric feature point statistics of the training data. The respective histograms display the total number of pharmacophoric feature points and the number of points of specific types per pharmacophore in the training data. The complete training dataset contains 1,217,361 distinct pharmacophores.
@@ A.2 Augmentation module
There are two modes for the displacement of pharmacophoric feature points, displacement within the tolerance sphere, and displacement onto the surface of the tolerance sphere. For simplicity, we assumed the same tolerance sphere radius $r_{T}$ across different pharmacophoric feature types. For displacement within the tolerance sphere, we calculated the coordinate displacement $(\Delta x,\Delta y,\Delta z)$ from spherical coordinates $\phi\sim\mathcal{U}(0,2\pi)$ and $\cos{\theta}\sim\mathcal{U}(-1,1)$, which were drawn at random from a uniform distribution:
@@
where $\Delta r=r_{T}\sqrt[3]{u}$ and $u\sim\mathcal{U}(0,1)$. Displacement of the nodes onto the tolerance sphere surface was achieved by calculating the mean of the positions of the pharmacophoric feature points, $\mathbf{\mu}=\frac{1}{|P|}\sum_{i=1}^{|P|}\mathbf{r}_{i}$, and displacing each point $p_{i}$ by $r_{T}$ in the direction $\mathbf{r}_{i}-\mathbf{\mu}$. The displacement away from the center ensures that displacement directions do not cancel each other.
@@ A.3 Message passing neural network
where $\mathbf{h}_{i}^{(k)}\in\mathbb{R}^{F^{\prime}}$ denotes the node features of node $i$ at layer $k$, $\mathbf{h}_{i}^{(k-1)}\in\mathbb{R}^{F}$ denotes the node features of node $i$ at layer $k-1$, $\mathbf{e}_{ij}\in\mathbb{R}^{D}$ the edge features of the edge from node $i$ to node $j$, $\gamma^{(k)}$ and $\phi^{(k)}$ are parameterized, differentiable functions, and $\bigoplus$ is an aggregation operator like, e. g., the summation operator (Fey & Lenssen, 2019). In our encoder architecture, we employed the following edge-conditioned convolution operator, which was proposed both by Gilmer et al. (2017) and Simonovsky & Komodakis (2017):
@@
where $\Theta\in\mathbb{R}^{F\times F{{}^{\prime}}}$ denotes learnable weights and $\psi_{\Theta}(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}^{F\times F{{}^{\prime}}}$ denotes a neural network, in our case an MLP with one hidden layer. These transformations map node features $\mathbf{h}$ into a latent representation that combines pharmacophoric feature types with distance encodings.
@@ A.4 Encoder implementation
where $\mathbf{h}^{l^{\prime}}_{i}$ represents the latent representation after activation. Updating the feature matrix $l$ times yields the final node representations of the pharmacophoric feature points.
@@
Figure 6: Architecture of the GNN encoder model.
@@ A.5 Model implementation and training
The GNN was implemented in Python 3.10 with PyTorch (v2.0.1) and the PyTorch Geometric library (v2.3.1) (Fey & Lenssen, 2019). Both, model and dataset, were implemented within the PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) framework (v2.1.0). Model training was monitored with Tensorboard (v2.13.0). CDPKit (v1.1.1) was employed for chemical data processing. Software was installed and executed on a Rocky Linux (v9.4) system with x86-64 architecture.
@@
Training was performed on a single NVIDIA GeForce 3090 RTX graphics unit with 24 GB GDDR6X. Training runs were performed for a maximum of 500 epochs with a batch size of 256 pharmacophore graphs. Curriculum learning was applied by gradual enrichment of the dataset with increasingly larger pharmacophore graphs. At training start, only pharmacophore graphs with 4 nodes were considered. After 10 subsequent epochs without considerable minimization of the loss function, pharmacophore graphs with one additional node were added to the training data. This approach allows the model to start with very simple examples, gradually increasing the difficulty of the matching task. The loss function was minimized with the Adam (Kingma, 2014) optimizer, we further applied gradient clipping. A training run on the full dataset took approximately 48 hours with the above hardware specifications.
Hyperparameter optimization & model selection
Hyperparameters were optimized through random parameter selection, the tested ranges are summarized in Table 4. Unlabeled data was split into training and validation data with a 98:2 ratio. Training runs were compared using the AUROC value on the validation data. This was calculated by treating the positive and negative pairs as binary labels, and the predictions were based on their respective order embedding penalty, which was calculated with function $E(\cdot,\cdot)$ in Equation (1). Hyperparameter optimization was performed on a reduced dataset with 100,000 graphs, which took approximately 5 hours per run. The best performing models were retrained on the full dataset. The hyperparameters of the final encoder model are summarized in Table 3. After model selection, the final model performance was tested on virtual screening datasets.
Table 3: Hyperparameters of the best performing encoder model
@@
Table 4: Tested hyperparameter ranges for model training.
Hyperparameter Parameter range
@@
Table 5: Tested hyperparameter ranges for ablation studies.
Parameter Parameter range Validation AUROC
Tolerance radius [0.0, 0.5, 1.0, 1.5, 2.0] [0.91, 0.93, 0.94, 0.94, 0.93]
Encoder dimension [8, 16, 32, 64, 96] [0.92, 0.93, 0.94, 0.94, 0.94]
Embedding dimension [32, 64, 128, 256, 512, 1024] [0.91, 0.93, 0.94, 0.94, 0.94, 0.93]
Skip-connection [dense, res, none] [0.94, 0.93, 0.74]
GNN Layer [NNConv, GINE, CFConv, GAT] [0.94, 0.94, 0.94, 0.93]
Projector layers [1, 2, 3] [0.94, 0.94, 0.94]
Convolution layers [1, 2, 3] [0.94, 0.94, 0.94]
Margin [0.01, 0.1, 1, 2, 5, 10, 100, 1000] [0.88, 0.90, 0.92, 0.92, 0.93, 0.93, 0.94, 0.94]
Ablation studies
To evaluate the importance of various model parameters, we conduct a series of ablation studies using the best-performing model. In these experiments, we systematically alter one parameter at a time and assess its impact on classification performance using the validation hold-out set. Our findings reveal several key insights. The embedding dimension of the learned representations can be reduced to 128 without loss in performance. The encoder requires at least 32 dimensions to remain effective. Skip-connections are critical for model performance, with DenseNet-style connections slightly outperforming ResNet-style (He et al., 2015) alternatives. Interestingly, the choice of message-passing layer, whether NNConv, graph isomorphism operator (GINE) (Hu et al., 2020), graph attention operator (GAT) (Brody et al., 2022), or continuous-filter convolutional layers (CFConv) (Schütt et al., 2018), has minimal impact on performance. The depth of the projector and encoder also does not significantly affect results. In contrast, the margin value plays a significant role in model performance. While larger values enhance performance, excessively high margins can lead to training instability. A margin of 100 provides an effective balance between these factors. The displacement radius for augmentations in creating positive pairs is most effective at 1.5 Å, while removing node displacement degrades model performance.
@@ A.6 Virtual screening
General information about the DUD-E targets is summarized in Table 6. For each target we downloaded the receptor structure from the PDB and created the corresponding interaction pharmacophore with the CDPKit. Vector features were converted into undirected pharmacophoric feature points. The resulting pharmacophore queries (Figure 7) were used in our virtual screening experiments.
Table 6: DUD-E targets that were selected for bechmarking experiments in this study.
@@
Figure 7: Structure-based pharmacophore queries of ten targets of the DUD-E benchmark dataset.
DEKOIS2.0 dataset details
General information about the DEKOIS2.0 targets (Bauer et al., 2013). Each target is associated with a PDB four-letter code and a corresponding ligand ID. For each target, we downloaded the receptor structure and its respective ligand from the PDB and generated the interaction pharmacophore using the CDPKit. These queries were used without further refinement in our pre-screening experiments. For targets containing the small molecule ligand in multiple binding pockets, we randomly selected one pocket for pharmacophore generation. The SIRT2 target was excluded from our evaluation because its structure does not contain a small molecule ligand. Actives and decoys were processed analogous to the ligands in the DUD-E benchmark.
LIT-PCBA dataset details
For each target, the downloaded files include a SMILES file containing the active and inactive ligands, along with receptor structures provided as PDB files. To ensure a fair comparison with PharmacoNet, we created pharmacophore queries using the same receptor PDB files as in their study. These PDB files were selected based on a methodology described by (Shen et al., 2023) and are listed in Table S4 of the respective supporting information. Interaction pharmacophores were generated using CDPKit and used without further refinement in our pre-screening experiments. Active and inactive ligands were processed following the same protocol as for the DUD-E benchmark.
@@
where $S_{MFP}:\mathcal{P}\times\mathcal{P}\rightarrow\mathbb{N}$ counts the number of matched feature pairs and $S_{Geom}:\mathcal{P}\times\mathcal{P}\rightarrow[0,1)$ evaluates their geometric fit.
@@
We measured alignment runtimes using the psdscreen tool from the CDPKit with 128 threads on an AMD EPYC 7713 64-Core Processor, while embedding and matching runtimes with PharmacoMatch were recorded using an NVIDIA GeForce RTX 3090 GPU with 24 GB GDDR6X. We additionally used 8 CPU worker nodes for the embedding step. Runtime per pharmacophore was estimated by dividing the total runtime by the number of pharmacophores in each dataset, with the final estimate taken as the mean of ten runs. The results report the mean and standard deviation of these estimates across all ten datasets.
Performance metrics
The enrichment factor (EF) is formally defined as:
$EF_{\alpha}=\frac{n_{\alpha}}{n\cdot\alpha}$ (13)
with $n_{\alpha}$ the number of true active compunds in the top $\alpha\%$ of the hitlist ranking and the total number of active compounds $n$.
The Boltzmann-Enhanced Discrimination of Receiver Operating Characteristic (BEDROC) metric (Truchon & Bayly, 2007) is similar to the AUROC metric but assigns greater weight to samples ranked higher in the hitlist. The degree of this reweighting is controlled by the early recognition parameter $\alpha$. The formal definition is as follows:
$BEDROC_{\alpha}={\sum_{i=1}^{n}e^{-\alpha r_{i}/N}\over{R_{a}({1-e^{-\alpha}\over{e^{\alpha/N}-1}})}}\cdot{R_{a}\sinh{(\alpha/2)}\over{\cosh{(\alpha/2)}-\cosh{(\alpha/2-\alpha R_{a})}}}+{1\over{1-e^{\alpha(1-R_{a})}}}$ (14)
where $n$ is the number of active compounds in the dataset, $N$ is the total number of compounds, $R_{a}=n/N$ is the ratio of active compounds in the dataset, and $r_{i}=\frac{\mathrm{rank}_{i}-1}{N-1}$ is the normalized ranking of the $i$-th active compound.
The area under receiver operating characteristic (AUROC) performance metrics of our virtual screening experiments on the selected DUD-E targets are derived from the ROC curves presented in Figures 8 and Figure 9. The performance metrics for the pre-screening experiments are calculated as follows: for each target, we calculated the AUROC, BEDROC, and enrichment factors (EF) for the top $0.5\%$, $1\%$, and $5\%$ of the hit list, reporting the average values for these metrics. Metrics for PharmacoNet were reported as presented in their original manuscript and were not recalculated.
@@ A.7 Embedding space visualization
Figure 8: Performance comparison of PharmacoMatch and the alignment algorithm. The ROC-curves display the agreement of the hitlist ranking of the two algorithms for ten targets of the DUD-E benchmark dataset.
Figure 9: Absolute screening performance of PharmacoMatch and the alignment algorithm performance for ten targets of the DUD-E benchmark dataset. The pharmacophore queries were generated from the respective PDB ligand-receptor structures.
Figure 10: UMAP visualization of the vector embeddings of the ACES target.
Figure 11: UMAP visualization of the vector embeddings of the ANDR target.
Figure 12: UMAP visualization of the vector embeddings of the EGFR target.
Figure 13: UMAP visualization of the vector embeddings of the FA10 target.
Figure 14: UMAP visualization of the vector embeddings of the KIT target.
Figure 15: UMAP visualization of the vector embeddings of the PLK1 target.
Figure 16: UMAP visualization of the vector embeddings of the SRC target.
Figure 17: UMAP visualization of the vector embeddings of the THRB target.
Figure 18: UMAP visualization of the vector embeddings of the UROK target.